# results on dynamics of bungee set of composite entire functions in the eremenko-lyubich class

### v2

@@ 3. Some results for functions of class $\mathcal{B}$
Let us consider a point $z_{0}\in RP(f)$. Then, $f^{n}(z_{0})$ is bounded and contained in $RP(f)$. Also, $g^{n}(RP(f))\subset RP(f)$, therefore $f^{n}(g^{n}(f^{n}(z_{0})))$ is also bounded as $n\rightarrow\infty.$ (Observe that, $z_{0}$ can’t be in $F(f\circ g)$ as $F(f\circ g)=F(f)$). Hence, $z_{0}\in RP(f\circ g)$ as well. On similar lines, if $w_{0}\in RP(g)$, then $w_{0}\in RP(f\circ g)$.
Now the question is can there be points not in $RP(f)\cup RP(g)$ but in $RP(f\circ g)$? To answer that let us go through (3.4) problem once again. When suppose $g(BU(f))$ goes to $RP(f)$ and upon further iteration with $(g\circ f)^{n}(z_{0})$, this will be bounded means $z_{0}\in RP(f\circ g)$. Similarly, there will be some points from $I(f)$ that will be bounded upon $g^{n}\circ f^{n}(z_{0})$ as $n\rightarrow\infty$ as discussed in (3.5). Again it has to be in $BU(g)$ and $I(g)$. So, this implies $RP(f\circ g)\supset RP(f)\cup RP(g)$. ∎
There’s a simpler proof for this which goes as follows:
@@
(3.3) $\implies RP(f\circ g)\supset RP(f)\cup RP(g)$
@@ 4. Illustrations:
We now illustrate $BU(f\circ g)\subset BU(f)\cup BU(g)$ with an example.
@@
Now, we discuss exponential functions which are not completely exponential type. Some examples are (see [15]): let us take the case where $f=p(z)\cdot e^{g(z)}$ where $p(z),\ g(z)$ are polynomials and $g(z)=a\cdot f(z)$ for some $a\in\mathbb{C}$. $f(z)=z\cdot e^{a^{(k+1)}\cdot z^{k}},\ k\in\mathbb{N}$ $g(z)=a\cdot z\cdot e^{a^{(k+1)}\cdot z^{k}}$.
@@
$g\circ f(z)=g(f(z))=a\cdot f(z)\cdot e^{a^{k}\cdot f^{k}(z)}=aze^{a^{k}z^{k}}e^{a^{2k}z^{k}e^{ka^{k}z^{k}}}.$ Now, for $f\circ g=g\circ f,$ we need to have $a^{2k+1}=a^{k}\implies a^{k}(a^{k+1}-1)=0$ which means $a$ is $(k+1)$-root of unity except ‘1’ itself because that’s trivial.
Now, for the simplest case let us take: $k+1=2\implies a=\pm 1$ Since a=1 is a trivial case so let us take $a=-1$, $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$. Here as the orbit of any point under either of the function tends to $\infty,f(z)\rightarrow 0,g(z)\rightarrow 0$ and $z=0$ is a fixed point of $f(z),\ g(z)$ since $f(0)=0$ and $g(0)=0$. Now for the multiplier $\lambda$, $f^{\prime}(z)=e^{-z^{2}}(1-2z^{2}),f^{\prime}(0)=1$ which is an indifferent fixed point. And $11^{1}=$ 1 means $0\in J(f)$ or $J(g)$ (since both of them are equal). As $\infty$ is a special case and belongs to $J(f)$ so, this proves that a sequence which tends to $\infty$ if it asymptotes to a finite asymptotic value which is a member of the Julia set of $f$. Moreover, the function which are of the form $f=p(z)e^{g(z)}$ which asymptotes because it has exponential term, means the asymptotic value is ${}^{\prime}0^{\prime}$ and it is in $J(f)$ (rationally-indifferent fixed point).
@@
Now, for the general case: Since $a$ is $k$-th root of unity, means $a^{k}=1$. So, $f(z)=ze^{a^{k}z^{k}}=ze^{z^{k}}$ and $g(z)=aze^{z^{k}}$. Now $g^{k}(z)=a^{k}\cdot f^{k}(z)=f^{k}(z)\implies BU(f)=BU(g)$. Also, $(f\circ g)^{k}(z)=a^{k}f^{2k}(z)=f^{2k}(z)$. So, $BU(f\circ g)=BU(f)=BU(g)$. Analogously, if we take $f(z)=ze^{a^{(k+1)}z^{k}}$ and $\ g(z)=af(z)=aze^{a^{(k+1)}z^{k}}$, where again $a^{k+1}=1$, we can simplify this further as $f(z)=ze^{az^{k}}$ and $g(z)=aze^{az^{k}}$, now $g^{2}(z)=af(af(z)))=a^{2}f^{2}(z)$. So, like this $g^{k+1}(z)=a^{(k+1)}f^{k+1}(z)$. Also $f\circ g(z)=af^{2}(z)=g\circ f(z)$. Now, if we do $(f\circ g)^{2}(z)=a^{2}f^{4}(z)$. So, $(f\circ g)^{k+1}(z)=a^{(k+1)}f^{2(k+1)}(z)=f^{2(k+1)}(z)$. Here also we can see that $BU(f\circ g)=BU(f)=BU(g)$.
@@ 5. Conclusion:
We have shown that entire functions with no finite asymptotic values which are also commuting i.e, $f\circ g=g\circ f$ satisfies:
$K(f)=K(g),\ I(f)=I(g),\ BU(f)=BU(g)$ and therefore $K(f\circ g)=K(f),\ BU(f\circ g)=BU(f),\ I(f\circ g)=I(f)$.
@@
There are some other special cases where we have equality among them like when $f$ is an exponential family and cases like $g=f+C$ where $C\in\mathbb{C}$ which are periodic translate of $f$. Also, when functions are from class $\mathcal{B}$, $(BU(f)\cup BU(g))\cap RP(f\circ g)\neq\emptyset$ because $g(BU(f))\subset RP(f)\cup BU(f)$. Similarly, $(I(f)\cup I(g))\cap RP(f\circ g)\neq\emptyset$ because $g(I(f))\subset RP(f)\cup I(f)$. Also, there are no points which are in $I(f)\cup I(g)$ as well as in $BU(f\circ g)$.

### v3

@@ 3. Some results for functions of class $\mathcal{B}$
(3.3) $\implies RP(f\circ g)\supset RP(f)\cap RP(g)$
@@ 4. Illustrations:
We now illustrate $BU(f\circ g)\subset BU(f)\cup BU(g)$ with some examples. Our first example is commuting rational functions.
@@
Our next example is motivated from [15]. Let us take the case where $f=p(z)\cdot e^{h(z)}$ where $p(z),\ h(z)$ are polynomials and $g(z)=a\cdot f(z)$ for some $a\in\mathbb{C}\setminus\{0\}$. For simplicity, let us take $f(z)=z\cdot e^{a^{(k+1)}\cdot z^{k}},\ k\in\mathbb{N}$ $g(z)=a\cdot z\cdot e^{a^{(k+1)}\cdot z^{k}}$.
@@
$g\circ f(z)=g(f(z))=a\cdot f(z)\cdot e^{a^{k}\cdot f^{k}(z)}=aze^{a^{k}z^{k}}e^{a^{2k}z^{k}e^{ka^{k}z^{k}}}.$ Now, for $f\circ g=g\circ f,$ we need to have $a^{2k+1}=a^{k}\implies a^{k}(a^{k+1}-1)=0,$ which means $a$ is $(k+1)$-root of unity except ‘1’ itself because that’s trivial.
Now, for the simplest case let us take: $k+1=2\implies a=\pm 1$ Since a=1 is a trivial case so let us take $a=-1$, $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$. Here as the orbit of any point under either of the function tends to $\infty,f(z)\rightarrow 0,g(z)\rightarrow 0$ and $z=0$ is a fixed point of $f(z),\ g(z)$ since $f(0)=0$ and $g(0)=0$. Now for the multiplier $\lambda$, $f^{\prime}(z)=e^{-z^{2}}(1-2z^{2}),f^{\prime}(0)=1$ which is an indifferent fixed point. And $11^{1}=$ 1 means $0\in J(f)$ or $J(g)$ (since both of them are equal). As $\infty$ is a special case and belongs to $J(f)$ so, this proves that a sequence which tends to $\infty$ if it asymptotes to a finite asymptotic value which is a member of the Julia set of $f$. Moreover, the functions which are of the form $f=p(z)e^{h(z)}$ have a finite asymptotic value ${}^{\prime}0^{\prime}$ and it is in $J(f)$ (rationally indifferent fixed point or repelling periodic point).
@@
Now, for the general case: Since $a$ is $k$-th root of unity, means $a^{k}=1$. So, $f(z)=ze^{a^{k}z^{k}}=ze^{z^{k}}$ and $g(z)=aze^{z^{k}}$. Now $g^{k}(z)=a^{k}\cdot f^{k}(z)=f^{k}(z)\implies BU(f)=BU(g)$. Also, $(f\circ g)^{k}(z)=a^{k}f^{2k}(z)=f^{2k}(z)$. So, $BU(f\circ g)=BU(f)=BU(g)$. Analogously, if we take $f(z)=ze^{a^{(k+1)}z^{k}}$ and $\ g(z)=af(z)=aze^{a^{(k+1)}z^{k}}$, where again $a^{k+1}=1$, we can simplify it further as $f(z)=ze^{az^{k}}$ and $g(z)=aze^{az^{k}}$, now $g^{2}(z)=af(af(z)))=a^{2}f^{2}(z)$. By induction, $g^{k+1}(z)=a^{(k+1)}f^{k+1}(z)$. Also, $f\circ g(z)=af^{2}(z)=g\circ f(z)$ and $(f\circ g)^{2}(z)=a^{2}f^{4}(z)$. Proceeding by induction, $(f\circ g)^{k+1}(z)=a^{(k+1)}f^{2(k+1)}(z)=f^{2(k+1)}(z)$. Here also, we can see that $BU(f\circ g)=BU(f)=BU(g)$.
@@ 5. Conclusion:
We have shown that entire functions with no finite asymptotic values which are also commuting i.e., $f\circ g=g\circ f$ satisfies:
$K(f)=K(g),\ I(f)=I(g),\ BU(f)=BU(g)$ and therefore, $K(f\circ g)=K(f),\ BU(f\circ g)=BU(f),\ I(f\circ g)=I(f)$.
@@
There are some other special cases where we have equality among these sets, like when $f$ is an exponential map. Also, cases like $g=f+C$ where $C\in\mathbb{C}$ is the period of $f$. Also, when functions are from class $\mathcal{B}$, $(BU(f)\cup BU(g))\cap RP(f\circ g)\neq\emptyset$ because $g(BU(f))\subset RP(f)\cup BU(f)$. Similarly, $(I(f)\cup I(g))\cap RP(f\circ g)\neq\emptyset$ because $g(I(f))\subset RP(f)\cup I(f)$. Also, there are no points which are in $I(f)\cup I(g)$ as well as in $BU(f\circ g)$.